13. Suppose T, S : R^2"\to" R^2 are linear defined by T(u, v) =(3u + v, u + 2v) and S(x, y) =(2x - y, x + y). Also the matrices of T and S with respect to the standard bases of R^2 and R^2 are given as

M(T) ="\begin{bmatrix}
 3 & 1 \\
 1 & 2
\end{bmatrix}" and M(S) ="\begin{bmatrix}
 2 & -1 \\
 1 & 1
\end{bmatrix}"Then M(TS) =

(i)"\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}"

(ii)"\begin{bmatrix}
 3 & 1 \\
 1 & 5
\end{bmatrix}"

(iii)"\begin{bmatrix}
 3 & 1 \\
 4 & 1
\end{bmatrix}"

(iv) None

14. Suppose T : R^2"\to" R^2 is linear defined by T(x, y) = (y, x). Then the eigenvalues of T is...

(i) 1 and - 1

(ii) 0 and 2

(iii) Does not exist

(iv) None


1
Expert's answer
2022-01-09T13:32:59-0500

13)


"M(TS)=\begin{bmatrix}
 3 & 1 \\
 1 & 2
\end{bmatrix}\begin{bmatrix}
 2 & -1\\
 1 & 1
\end{bmatrix}=\begin{bmatrix}
 7 & -2 \\
 4 & 1
\end{bmatrix}"


Correct option: None


14)


"T(x,y)=(y,x)"


"\implies\>T(x,y)=\lambda(x,y)"

"\therefore\>\lambda(x,y)=(y,x)"


"\lambda\>x=y........(i)"

"\lambda\>y=x........(ii)"




Substituting "(ii)" in "(i)"


"\lambda(\lambda\>y)=y"

"\lambda^2=1"

"\lambda=^+_-1"


"\therefore\lambda=1\>or\>-1"